the kinetic energy of an electron in H like atom is 6.04 eV. the area of the third bohr orbit to which this electron belong is;

Grade:11

1 Answers

Arun
6 years ago
K.E. = 6.04 ev = 9.7 * 10^-19 J
K.E. of electron = Kze²/2r
K = constant = 9* 10^9 Nm²/C²
z = 1( for hydrogen like atom)
e = 1.6 * 10^-19 C
Hence put all this in the above formula
You will get
r = 1.19 * 10^-10 metres
Hence area A = (3.14159) * r²
A= 17.8 * 10^-20 m²